OpenStudy (fairy733189):

The table below shows the sale, in dollars, at Jacob's store over a period of five months: Month: 1 2 3 4 5 Sale:1,000 1,050 1102.50 1157.63 1215.51 Did the number of people at Jacob's store increase linearly or exponentially? (4 points) Linearly, because the table shows a constant percentage increase in sales per month Exponentially, because the table shows a constant percentage increase in sales per month Linearly, because the table shows that sales increase by an equal factor for an equal increase in months Exponentially because the table shows and equal increase in sales for an equal increase in months

OpenStudy (reemii):

What is the percentage increase from Month1 to Month2?

OpenStudy (fairy733189):

I'm sorry but I'm terrible at math so I don't really know How do you find that out?

OpenStudy (fairy733189):

I actually want to learn

OpenStudy (fairy733189):

is it 50?

OpenStudy (reemii):

Value \(x\) becomes \(y\). Divide \(y\) by \(x\): \(\frac{1050}{1000} = 1.05\) The increase if \(1.05 - 1\). (always subtract \(1\)). For example, 1000->2000 means a \(\frac{2000}{1000} - 1 = 2-1=1 = \frac{100}{100} = 100\%\) increase.

OpenStudy (reemii):

*if = is

OpenStudy (fairy733189):

so would it be 5%?

OpenStudy (reemii):

Yes

OpenStudy (fairy733189):

Great!! So what do i do next?

OpenStudy (reemii):

You can compute all %increases, ......... it will always be 5%. (trust me). So we are in front of a "constant % increase of sales" (not an increase in percentage!)

OpenStudy (fairy733189):

ok but one question that I have is what is the difference between a constant increase and an exponential increase?

OpenStudy (reemii):

(without searching…) from memory, LINEAR means: "two months pass, I win twice what I win in a month, whatever 2 months we are looking at." Not the case here.

OpenStudy (reemii):

Exponential means: start anywhere with some value \(A\). After \(k\) months, the value is \(A(1+r)^k\) (where percentage is \(r\)). That is the case here. You (or I) have checked that between any two value, the %increase is 5%. That's what had to be checked.

OpenStudy (fairy733189):

So linear is when the percentage doesn't change and exponential is when it does over a certain amount of time?

OpenStudy (reemii):

Not really. Let's call the values \(s_1, s_2,\dots,s_5\). We saw that the increase from \(s_1\) to \(s_2\) is 5%. If it were linear, the increase from Month 1 to Month 3 would be \(+(2\times 5\%)\). But, here, \(s_1\to s_2\) (+5%), \(s_2\to s_3\) (+5%) etc. \(s_4\to s_5\) (+5%) Since an increase of 5% from value X to Y can be written: Y = X (1+ 0.05), our sales verify this: \[s_4 = s_3(1+0.05) = s_2 (1+0.05)(1+0.05) \\= s_1(1+0.05)(1+0.05)(1+0.05)=s_1(1+0.05)^3\]. It is exponential.

OpenStudy (fairy733189):

oh ok

OpenStudy (fairy733189):

So after finding this what do we do next?

OpenStudy (reemii):

An increase in \(r\%\) (written as a number between 0 and 1) means multiplying by a factor\((1+r\).

OpenStudy (reemii):

\()\).

OpenStudy (reemii):

There are two lines starting with "Exponentially". One is wrong.

OpenStudy (fairy733189):

But one says constant and the other says equal

OpenStudy (reemii):

it's the following noun that counts: "constant percentage increase" or "constant increase in sales" ?

OpenStudy (fairy733189):

so it would be the second option

OpenStudy (fairy733189):

out of all 4

OpenStudy (reemii):

I think so too.

OpenStudy (fairy733189):

Awesome!! Thank You!!!

OpenStudy (reemii):

yw Fairy!